OpenStudy (anonymous):

A square-shaped deck has an area of 289 ft2. What is the perimeter of the deck? ____________ft

OpenStudy (anonymous):

Area = base * height

OpenStudy (anonymous):

you have the area and you know its a square

OpenStudy (anonymous):

so Area = x * x

OpenStudy (anonymous):

Okay so is it simple to find the perimeter

OpenStudy (anonymous):

what is 289 = x^2 solve x

OpenStudy (anonymous):

uhhh I'm confused at this bit

OpenStudy (anonymous):

1.01?

Mehek (mehek14):

find the \(\Large \sqrt{289}\)

OpenStudy (anonymous):

17?

Mehek (mehek14):

ok so that's the length of one side

Mehek (mehek14):

for perimeter just multiply 17 x 4

OpenStudy (anonymous):

68?

Mehek (mehek14):

yup

Mehek (mehek14):

so that's the perimeter

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

So it is actually 68?

Mehek (mehek14):

yea...